32 over 27 equals x over 15. do u no the answer
13 over n equals 10 over 6. do u no the answer
a map has a scale of 7 cm represisents 8 km. which proportion would not allow u to find the actual distance between two cities if they are 9 cm apart on the map?
which of the following fractions has the greatest value.
22over21 20over13 19over20 28over37
by the way i need to explain how i got all of these answers

Answer by uknouluvme89(30)

set up a proportion
32/27 = x/15
cross multiply
480 = 27x
divide both sides by 27
17.77777778 = x (round it to whatever decimal place you need)
13/n = 10/6
78=10x
7.8 = x